# Scalable Stewardship of an LLM-Assisted Clinical Benchmark with Physician Oversight

### v2

@@ Appendix E Controlled RL Experiment Details
Both training runs are conducted on an $8\timesH$100 GPU node with two 56-core Intel 8480C CPUs for 9 hours, using the verl framework for distributed RL training.[35] The experiment run in Fig. 3 consumes $2\times(8\times 9)=144$ GPU hours; if including the sensitivity analysis, our experiment consumes $144\times(3\times 3)=1296$ GPU hours.
@@ Appendix F Two $\mathsf{N/A}$ Instances Found by Phase 1 Audit
“A 58-year-old male with dyslipidemia, an eight-year history of T2DM, a family history, his mother, of T2DM, with no known micro- or macrovascular complications, was admitted to the emergency department for malaise, epigastric pain, polyuria, and progressive dyspnea which had begun 10 h ago. He had experienced a 2-kg weight loss over the last few days. His usual medications included aspirin 100 mg q24 h, atorvastatin 40 mg q24 h, and metformin 850 mg q8 h, which had been switched to dapagliflozin 20 days before, due to poor glycemic control, with HbA1c 12% (108 mmol/mol). His vital signs included a heart rate of 122 bpm, respiratory rate 33 rpm, blood pressure 142/70 mmHg, temperature 36.1°C, and body mass index 22.5 kg/m2. On physical examination, somnolence, dry skin and mucous membranes, a Kussmaul breathing pattern, and a capillary refill of 3 sec were observed. Blood tests revealed hemoglobin 17.1 g/dL (13.5–18), leukocytes 19.5 × 103 (4–10 × 103), platelets 296 × 103 (150–450 × 103), glucose 248 mg/dL (60–100), creatinine 0.97 mg/dL (0.67–1.17), sodium 136 mmol/L (135–145), potassium 4.7 mmol/L (3.5–5.5), chloride 101 mmol/L (95–112), phosphate 4.9 mg/dL (2.5–4.5), amylase 70 U/L (10–115), lipase 28 U/L (1–67), pH 6.95 (7.35–7.45), pCO2 23 mmHg (35–45), HCO3 5 mmol/L (22–26), lactate 1.8 mmol/L (0–1.5), urine ketone bodies >150 mg/dL (0–0), CK 112 U/L (1–190), CK-MB 7.3 ng/mL (0.1–5), and troponin I 0.07 ng/mL (0.001–0.05). The electrocardiogram (EKG) showed sinus rhythm with right bundle branch block, and nonspecific repolarization abnormalities. Because of the right bundle branch block was not previously known, a new troponin test was performed six hours later with a peak value of 4.28 ng/mL. Treatment with crystalloids, continuous infusion of intravenous insulin, and administration of potassium and sodium bicarbonate were begun in the emergency room (ER). Due to a poor response over the first two hours, with the persistence of lactic acidosis, the patient was transferred to the intensive care unit (ICU), where more aggressive rehydration with crystalloids was started, without further modifications of the original therapeutic plan. Two days later, the patient was discharged from the ICU to the endocrinology ward. Because of his coronary risk factors and the elevated troponin on admission, a coronary angiography was performed, showing triple-vessel disease. Successful bypass surgery without extracorporeal circulation was performed a few days later, with internal mammary artery grafts to the anterior descendent and marginal obtuse arteries and a saphenous vein graft to the right coronary artery. He was discharged 3 days later on Lantus™ Sanofi-aventis S.p.a Valcanello,03012 Anagni (FR), Italia (insulin glargine) 20 IU and Insulina Novorapid™: Novo Nordisk A/S. Hallas Allé, DK-4400. Kalundborg, Dinamarca (insulin aspart) 6-4-4-0 IU subcutaneous insulin, aspirin, clopidogrel, enalapril, bisoprolol, atorvastatin, and furosemide. The oral antidiabetic treatment with dapagliflozin was not restarted. During follow-up by endocrinology, C-peptide, anti-GAD, and IA-2 antibodies were required due to normal BMI and lack of family history suggested T1DM or a latent autoimmune diabetes of adults (LADA) rather than T2DM as etiology of his diabetes. C-peptide was 1 mg/dL (0.9–7.1) with glycemia 214 mg/dL (60–100), anti-GAD <5 U/mL (0–12.5), and anti-IA-2 <7.5 U/mL (0–7.5).”
@@
“A 55-year-old woman presented with a past medical history of smoking until five years ago (60 pack/year), arterial hypertension, chronic obstructive pulmonary disease with no supplementary oxygen use, and chronic hepatitis B infection with no history suggestive of cirrhosis. Her chief complaint was one year of abdominal discomfort associated with an epigastric mass and weight loss of 30 kg. Simple and contrasted abdominal computed tomography (CT) showed hepatomegaly with focal lesions, the largest being 15 cm in segment VIII, and satellite lesions of up to 5.5 cm in segment VII and 3.5 cm in segment III, with similar morphologic characteristics. Abdominal magnetic resonance imaging (MRI) showed a liver with no signs of chronic disease, and an infiltrative dominant lesion of 11 x 8 cm in segments IV and V, associated with smaller lesions in both lobules of 1.6 cm in segment III, and 0.9 cm in segments VII and VIII (Figure ). Subsequently, the patient presented aggressiveness and disorientation, where hypoglycemia was documented (33 mg/dL) and treated with dextrose 10% with complete symptom resolution. Physical examination revealed cachectic facies, and an abdominal mass of approximately 13 cm located in epigastrium and right hypochondrium, associated with hepatomegaly. Initial laboratory were: white blood count of 9900/uL, neutrophils 8000/uL, lymphocytes 1700/uL, monocytes 300/uL, red blood cells 5,540,000/uL, hemoglobin 16.6 g/dL, hematocrit 49.6%, mean corpuscular volume 89.6 fL, platelet count 357,000/uL, international normalised ratio (INR) 1.05 (normal range (NR) 0.8-1.4), partial thromboplastin time 33 seconds (NR 25-35), serum potassium 3.3 mmol/L, chloride 104 mmol/L (NR 98-107), sodium 141 mmol/L (NR 136-145), total bilirubin 1.01 mg/dL (NR 0.2-0.9), direct bilirubin 0.48 mg/dL (NR 0-0.3), alkaline phosphatase 404 U/L (NR 38-110), gamma-glutamyl transferase (GGT) 1017.6 U/L (NR 10-38), lactic dehydrogenase (LDH) 293 U/L (NR 120-246), alanine transaminase (ALT) 124 U/L (NR 9-52), aspartate transaminase (AST) 226 U/L (NR 14-36), serum creatinine 0.22 mg/dL (NR 0.5-0.8), serum alpha fetoprotein (AFP) 23184.1 ng/mL (NR 0.8-1), carcinoembryonic antigen 0 ng/mL, CA 19-9 antigen 26.86 U/mL (NR 0-37), cortisol 8.9 ug/dL (NR 5.27-22.45), blood ketones 1.0 mg/dL (NR < 3.5), serum glucose 17 mg/dL (NR 74-106), baseline insulin 0.40 uU/mL (NR 2.6-24.9), IGF-I < 15 ng/mL (NR 45-210), C peptide 0.0407 ng/mL (NR 1.1-4.4), IgM antibodies against hepatitis B core antigen non-reactive, antibodies against hepatitis B surface antigen < 3.10 mUI/mL (NR 0-10), antibodies against hepatitis C non-reactive, hepatitis B surface antigen (HBsAg) > 1000 (NR 0-1). Given our patient’s profile, a “big” IGF-II measurement was attempted, but the assay is not available in Colombia”……
@@ Appendix G MedCalc-Bench Usage in Anthropic Press Release
In Anthropic’s “Claude for Healthcare” press release,[3] referenced earlier in the main text’s Introduction section (§1), the company did not provide the evaluation source code or the version number of MedCalc-Bench dataset with which their models obtained the reported accuracies, i.e., whether it was the original peer-reviewed version published in NeurIPS 2024 (v1.0), or a recent GitHub/HuggingFace revision released by the same authors without a technical report on how their reference labels’ curation methodology changed (v1.2). [^6] Appendix §B provides more context on the distinction between the two MedCalc-Bench versions. Identifying the dataset version used by Anthropic is essential for gauging the clinical validity of their LLMs’ benchmark results. To balance inference cost with experiment relevance, we tried to independently evaluate Anthropic’s Haiku 4.5| and
@@ G.2 Our Experiment Results and Analysis
Figure 7: Benchmark performance of four Claude-series models was reported by Anthropic in Exhibit #1 of their Jan 11, 2026 press release.[3] Screenshot taken on Jan 17, 2026. We configured our API call pipeline to match the setup described in the first two bullet-point footnotes. MedAgentBench [15] is unrelated to our study.

### v3

@@ Appendix E Controlled RL Experiment Details
Both training runs are conducted on an $8\timesH$100 GPU node with two 56-core Intel 8480C CPUs for 9 hours, using the verl framework for distributed RL training (Sheng et al., 2025). The experiment run in Fig. 9 consumes $2\times(8\times 9)=144$ GPU hours; if including the sensitivity analysis, our experiment consumes $144\times(3\times 3)=1296$ GPU hours.
E.1 Full Training Dynamics on 887-Instance Test Set
The main text (§3.4) reports the RL experiment results on the 50 physician-labeled instances. Here we present the full training dynamics evaluated on the larger 887-instance held-out test set, graded against $\hat{y}^{\mathrm{new}}$. Averaging the final ten evaluation steps, the model trained on $\hat{y}^{\mathrm{new}}$ achieves $71.4\%$ accuracy versus $62.6\%$ for the model trained on $\hat{y}^{\mathrm{original}}$, an $8.7$ percentage-point gap (Figure 9). Sensitivity analysis over a $3\times 3$ hyperparameter sweep confirms this difference is directionally consistent across all settings.
Notably, while the label discrepancy between $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$ is roughly 27% (§3.2), the downstream RL performance gap is smaller ($8.7\%$), suggesting that the RL process displays some resilience to noisy rewards. This $8.7\%$ swing can also be read as an empirical measure of evaluation uncertainty induced by label choice: changing only the reference labels, while holding the data, model class, and optimization fixed, can move the headline conclusion by $8.7\%$. For context, this difference could eliminate or double the gaps between the top three reported alignment results on MedCalc-Bench (Table 2 in Appendix D.2).
Figure 9: Test accuracy dynamics on the 887-instance held-out set for Qwen3-8B trained via GRPO using the recomputed reference labels (green) versus the original MedCalc-Bench labels (grey). Shaded bands indicate $\pm 1\sigma$ smoothed over a 10-step window. The model trained on $\hat{y}^{\mathrm{new}}$ achieves a $8.7\%$ absolute gain in the final moving averages of test accuracy ($71.4\%$ vs. $62.6\%$).
E.2 Generalization to Additional Medical Evaluation Sets
In this section, we test whether the performance gains reported in §3.4 generalize beyond the MedCalc-Bench test split used in the main experiment. We evaluate both RL-trained checkpoints on three additional medical evaluation sets:
  • MedQA (Jin et al., 2021): a medical question-answering benchmark that tests a fundamentally different task (multiple-choice clinical reasoning rather than score computation).
  • MedCalc-v1.2: a recent revision of MedCalc-Bench posted online by an author of the original benchmark (HuggingFace link). Our examination suggests that approximately 70% of test instances and annotations in the original v1.0 release were replaced, making this a partially overlapping but substantially different version of the same task.
  • MedCalc-Bench-Verified: another revision (GitHub link, December 14 snapshot), also posted by an author of the original benchmark, with similarly extensive changes to the v1.0 release.
Regression setup.
We fit a linear probability model with the binary outcome Accuracy (1 if the model’s answer is correct, 0 otherwise) as the dependent variable. The treatment variable is the training label source, coded as an indicator for orig_train (the checkpoint trained on $\hat{y}^{\mathrm{original}}$), with the checkpoint trained on $\hat{y}^{\mathrm{new}}$ (new_train) as the reference category. We also include the base Qwen3-8B checkpoint before either RL run (untrained) as a third level. Control variables are dataset (MedQA, MedCalc-v1.2, MedCalc-Bench-Verified) and decoding configuration (temperature $T$, top_p, top_k), which absorb differences in task difficulty and sampling variance across settings.
Results.
The regression was fit on 150,048 sampled generations ((1,146+990+990)(1{,}146+990+990) prompts $\times$ 3 checkpoints $\times$ 2 decoding settings $\times$ 8 sampled trajectories per prompt). The estimated coefficient for orig_train is 0.0093-0.0093 (95% CI, 0.015-0.015 to 0.004-0.004; $p=0.001$), corresponding to a 0.93 percentage-point accuracy deficit relative to new_train after controlling for dataset and decoding configuration. Table 3 reports the unadjusted mean accuracies for every dataset-decoding cell.
Table 3: Mean accuracy (%) on additional evaluation sets, grouped by dataset, decoding configuration, and checkpoint. The three checkpoints are the model trained on $\hat{y}^{\mathrm{new}}$ (new_train), the model trained on $\hat{y}^{\mathrm{original}}$ (orig_train), and the base model before RL (untrained). Decoding tuples are reported as ($T$, top_p, top_k). The $\Delta$ column reports new_train minus orig_train in percentage points.
Dataset ($T$, top_p, top_k) new_train orig_train untrained $\Delta$
MedQA $(0.6,0.95,20)$ 79.2 78.2 75.7 +1.0
MedQA (1.0,1.0,1)(1.0,1.0,-1) 78.4 78.2 74.5 +0.2
MedCalc-v1.2 $(0.6,0.95,20)$ 64.1 63.0 30.1 +1.1
MedCalc-v1.2 (1.0,1.0,1)(1.0,1.0,-1) 63.1 62.9 29.2 +0.2
MedCalc-Bench-Verified $(0.6,0.95,20)$ 64.2 62.3 29.7 +1.9
MedCalc-Bench-Verified (1.0,1.0,1)(1.0,1.0,-1) 63.0 61.6 29.4 +1.4
In all six dataset-by-decoding combinations, new_train outperforms orig_train, with raw gains ranging from 0.2 to 1.9 percentage points. The effect is directionally consistent but substantially smaller than the +8.7+8.7 percentage-point gain observed on the in-distribution MedCalc-Bench test split (Figure 9). This attenuation is expected: the RL fine-tuning signal is most concentrated on MedCalc-style score computation, and a smaller model like Qwen3-8B has limited capacity to transfer fine-tuning gains to tasks beyond its training distribution. Larger models may exhibit stronger generalization of label-quality effects, but testing this hypothesis is beyond the scope of the current study. The key finding is that training on higher-quality labels does not degrade performance on related medical tasks and, if anything, yields a modest, statistically significant improvement.
@@ Appendix F Two $\mathsf{N/A}$ Instances Found by Phase 1 Audit
“A 58-year-old male with dyslipidemia, an eight-year history of T2DM, a family history, his mother, of T2DM, with no known micro- or macrovascular complications, was admitted to the emergency department for malaise, epigastric pain, polyuria, and progressive dyspnea which had begun 10 h ago. He had experienced a 2-kg weight loss over the last few days. His usual medications included aspirin 100 mg q24 h, atorvastatin 40 mg q24 h, and metformin 850 mg q8 h, which had been switched to dapagliflozin 20 days before, due to poor glycemic control, with HbA1c 12% (108 mmol/mol). His vital signs included a heart rate of 122 bpm, respiratory rate 33 rpm, blood pressure 142/70 mmHg, temperature 36.1°C, and body mass index 22.5 kg/m2. On physical examination, somnolence, dry skin and mucous membranes, a Kussmaul breathing pattern, and a capillary refill of 3 sec were observed. Blood tests revealed hemoglobin 17.1 g/dL (13.5–18), leukocytes 19.5 × 103 (4–10 × 103), platelets 296 × 103 (150–450 × 103), glucose 248 mg/dL (60–100), creatinine 0.97 mg/dL (0.67–1.17), sodium 136 mmol/L (135–145), potassium 4.7 mmol/L (3.5–5.5), chloride 101 mmol/L (95–112), phosphate 4.9 mg/dL (2.5–4.5), amylase 70 U/L (10–115), lipase 28 U/L (1–67), pH 6.95 (7.35–7.45), pCO2 23 mmHg (35–45), HCO3 5 mmol/L (22–26), lactate 1.8 mmol/L (0–1.5), urine ketone bodies >>150 mg/dL (0–0), CK 112 U/L (1–190), CK-MB 7.3 ng/mL (0.1–5), and troponin I 0.07 ng/mL (0.001–0.05). The electrocardiogram (EKG) showed sinus rhythm with right bundle branch block, and nonspecific repolarization abnormalities. Because of the right bundle branch block was not previously known, a new troponin test was performed six hours later with a peak value of 4.28 ng/mL. Treatment with crystalloids, continuous infusion of intravenous insulin, and administration of potassium and sodium bicarbonate were begun in the emergency room (ER). Due to a poor response over the first two hours, with the persistence of lactic acidosis, the patient was transferred to the intensive care unit (ICU), where more aggressive rehydration with crystalloids was started, without further modifications of the original therapeutic plan. Two days later, the patient was discharged from the ICU to the endocrinology ward. Because of his coronary risk factors and the elevated troponin on admission, a coronary angiography was performed, showing triple-vessel disease. Successful bypass surgery without extracorporeal circulation was performed a few days later, with internal mammary artery grafts to the anterior descendent and marginal obtuse arteries and a saphenous vein graft to the right coronary artery. He was discharged 3 days later on Lantus™ Sanofi-aventis S.p.a Valcanello,03012 Anagni (FR), Italia (insulin glargine) 20 IU and Insulina Novorapid™: Novo Nordisk A/S. Hallas Allé, DK-4400. Kalundborg, Dinamarca (insulin aspart) 6-4-4-0 IU subcutaneous insulin, aspirin, clopidogrel, enalapril, bisoprolol, atorvastatin, and furosemide. The oral antidiabetic treatment with dapagliflozin was not restarted. During follow-up by endocrinology, C-peptide, anti-GAD, and IA-2 antibodies were required due to normal BMI and lack of family history suggested T1DM or a latent autoimmune diabetes of adults (LADA) rather than T2DM as etiology of his diabetes. C-peptide was 1 mg/dL (0.9–7.1) with glycemia 214 mg/dL (60–100), anti-GAD <<5 U/mL (0–12.5), and anti-IA-2 <<7.5 U/mL (0–7.5).”
@@
“A 55-year-old woman presented with a past medical history of smoking until five years ago (60 pack/year), arterial hypertension, chronic obstructive pulmonary disease with no supplementary oxygen use, and chronic hepatitis B infection with no history suggestive of cirrhosis. Her chief complaint was one year of abdominal discomfort associated with an epigastric mass and weight loss of 30 kg. Simple and contrasted abdominal computed tomography (CT) showed hepatomegaly with focal lesions, the largest being 15 cm in segment VIII, and satellite lesions of up to 5.5 cm in segment VII and 3.5 cm in segment III, with similar morphologic characteristics. Abdominal magnetic resonance imaging (MRI) showed a liver with no signs of chronic disease, and an infiltrative dominant lesion of 11 x 8 cm in segments IV and V, associated with smaller lesions in both lobules of 1.6 cm in segment III, and 0.9 cm in segments VII and VIII (Figure ). Subsequently, the patient presented aggressiveness and disorientation, where hypoglycemia was documented (33 mg/dL) and treated with dextrose 10% with complete symptom resolution. Physical examination revealed cachectic facies, and an abdominal mass of approximately 13 cm located in epigastrium and right hypochondrium, associated with hepatomegaly. Initial laboratory were: white blood count of 9900/uL, neutrophils 8000/uL, lymphocytes 1700/uL, monocytes 300/uL, red blood cells 5,540,000/uL, hemoglobin 16.6 g/dL, hematocrit 49.6%, mean corpuscular volume 89.6 fL, platelet count 357,000/uL, international normalised ratio (INR) 1.05 (normal range (NR) 0.8-1.4), partial thromboplastin time 33 seconds (NR 25-35), serum potassium 3.3 mmol/L, chloride 104 mmol/L (NR 98-107), sodium 141 mmol/L (NR 136-145), total bilirubin 1.01 mg/dL (NR 0.2-0.9), direct bilirubin 0.48 mg/dL (NR 0-0.3), alkaline phosphatase 404 U/L (NR 38-110), gamma-glutamyl transferase (GGT) 1017.6 U/L (NR 10-38), lactic dehydrogenase (LDH) 293 U/L (NR 120-246), alanine transaminase (ALT) 124 U/L (NR 9-52), aspartate transaminase (AST) 226 U/L (NR 14-36), serum creatinine 0.22 mg/dL (NR 0.5-0.8), serum alpha fetoprotein (AFP) 23184.1 ng/mL (NR 0.8-1), carcinoembryonic antigen 0 ng/mL, CA 19-9 antigen 26.86 U/mL (NR 0-37), cortisol 8.9 ug/dL (NR 5.27-22.45), blood ketones 1.0 mg/dL (NR << 3.5), serum glucose 17 mg/dL (NR 74-106), baseline insulin 0.40 uU/mL (NR 2.6-24.9), IGF-I << 15 ng/mL (NR 45-210), C peptide 0.0407 ng/mL (NR 1.1-4.4), IgM antibodies against hepatitis B core antigen non-reactive, antibodies against hepatitis B surface antigen << 3.10 mUI/mL (NR 0-10), antibodies against hepatitis C non-reactive, hepatitis B surface antigen (HBsAg) >> 1000 (NR 0-1). Given our patient’s profile, a “big” IGF-II measurement was attempted, but the assay is not available in Colombia”……
@@ Appendix G MedCalc-Bench Usage in Anthropic Press Release
In Anthropic’s “Claude for Healthcare” press release (Anthropic, 2026a), referenced earlier in the main text, the company did not provide the evaluation source code or the version number of MedCalc-Bench dataset with which their models obtained the reported accuracies, i.e., whether it was the original peer-reviewed version published in NeurIPS 2024 (v1.0), or a recent GitHub/HuggingFace revision released by the same authors without a technical report on how their reference labels’ curation methodology changed (v1.2). [^6] Appendix §B provides more context on the distinction between the two MedCalc-Bench versions. Identifying the dataset version used by Anthropic is essential for gauging the clinical validity of their LLMs’ benchmark results. To balance inference cost with experiment relevance, we tried to independently evaluate Anthropic’s Haiku 4.5 and Sonnet 4.5 models using the Claude API, and then compare our results with their results shown in Figure 10.
@@ G.2 Our Experiment Results and Analysis
Figure 10: Benchmark performance of four Claude-series models was reported by Anthropic in Exhibit #1 of their Jan 11, 2026 press release (Anthropic, 2026a). Screenshot taken on Jan 17, 2026. We configured our API call pipeline to match the setup described in the first two bullet-point footnotes. MedAgentBench (Jiang et al., 2025) is unrelated to our study.